# Emerging Quadrature Lattices of Kerr Combs

### v1

@@ Abstract
A quadrature lattice is a coupled array of squeezed vacuum field quadratures that offers new avenues in shaping the quantum properties of multimode light [1, 2, 3]. Such lattices are described within the framework of non-Hermitian, non-dissipative physics and exhibit intriguing lattice phenomena such as lattice exceptional points, edge-states, entanglement and non-Hermitian skin effect, offering fundamentally new methods for controlling quantum fluctuations [1, 4]. Nonlinear resonators are suitable for studying multimode pair-generation processes and squeezing which are non-dissipative in $\chi^{(2)}$ and $\chi^{(3)}$ materials [5, 6, 7, 8, 9, 10, 11, 12], but observing non-Hermitian lattice phenomena in photonic quadrature lattices was not achieved. Remarkably, in dissipative Kerr microcombs [13], which have revolutionized photonic technology, such lattices emerge and govern the quantum noise that leads to comb formation. Thus, they offer a unique opportunity to realize quadrature lattices, and to study and manipulate multimode quantum noise which is essential for any quantum technology. Here, we experimentally study non-Hermitian lattice effects in photonic quadrature lattices for the first time. Our photonic quadrature lattices emerge at Kerr microcomb transitions, allowing us to observe fundamental connections between dispersion symmetry, frequency-dependent squeezed supermodes, and non-Hermitian lattice physics in an integrated setup. Our work unifies two major fields, quantum non-Hermitian physics and Kerr combs, and opens the door to utilizing dissipative Kerr combs to experimentally explore rich non-Hermitian physics in the quantum regime, engineer quantum light, and develop new tools to study the quantum noise and formation of Kerr combs.
@@ 1 Introduction
In recent years, there has been growing interest in photonic quadrature lattices and their quantum properties [1]. Unlike conventional photonic lattices, where the lattice sites are coherent field amplitudes, in photonic quadrature lattices, each discrete site consists of two optical quadratures that are coupled to other quadratures. Since the quadrature representation forms the natural basis for continuous variable (CV) [14] quantum states, photonic quadrature lattices enable the understanding of rich multimode systems from the perspective of squeezing and deterministic entanglement generation. Notably, Hermitian quadratic bosonic lattices translate to non-Hermitian lattices in quadrature space. Thus, the rich physics of non-Hermitian lattices, which includes a plethora of effects such as optical exceptional points [15], the non-Hermitian skin effect [16], and beyond, manifests in quadrature space with novel implications for squeezing and entanglement [16, 17, 2, 4].
In contrast to traditional non-Hermitian optical lattices, where the interplay between gain and loss is governed by photon absorption, quadrature lattice dynamics are governed by parametric processes (i.e., energy-conserving) such as nonlinear optical pair creation and annihilation. Thus, the realization of a bosonic quadrature lattice requires both multimode parametric gain and loss, and the coupling or transfer of bosons between lattice sites. In recent experiments, non-Hermitian lattice effects in quadrature lattices were observed in optomechanics [2, 18, 3] by coupling several parametrically pumped opto-mechanical resonator modes together. In photonics, a fruitful approach for observing non-Hermitian lattice physics is along the synthetic frequency dimension of a single optical cavity, in which the optical modes are coupled by modulation at the free spectral range (FSR) frequency [19, 20, 21]. Such lattices can also be realized in a nonlinear environment [22], and even in the non-dissipative quantum regime by coupling photons generated across multiple frequency modes by second-order pair production with electro-optic modulation [8]. However, observing quadrature lattices requires detecting the multimode quadrature states with a local oscillator [23, 10, 11, 24], and probing the non-Hermitian lattice effects, which has not been achieved so far.
Remarkably, Kerr combs [13] (including dissipative Kerr solitons [25] and soliton crystals [26]) generated by third-order nonlinear resonators, naturally provide both pair production [27, 28] and coupling between the frequency modes through the process of Bragg scattering [5, 29, 30, 31]. These properties make Kerr combs a unique system to study multimode quantum phenomena [7, 32] and squeezing [33]. Interestingly, these Kerr combs also function as a multi-frequency pump (in contrast to a single pump [9]) which produces transport and amplification of the quantum fluctuations in the resonator, giving rise to “below-threshold” phase-locked combs [7]. The complexity of these squeezed combs, and their quadratic bosonic Hamiltonian, which includes parametric gain and coupling terms, has only recently begun to be treated theoretically [34, 33]. However, to date, such below-threshold multimode quadratures have not been measured directly, and it was not clear how these combs are related to quadrature lattices and the non-Hermitian lattice phenomena that govern them. Additionally the major development of dissipative Kerr microcombs, which has become the focal point of vast research efforts, offers new possibilities in shaping quantum light, but also pose new limitations. Thus, studying the multimode quantum squeezing and anti-squeezing generated by them is essential for both utilizing their quantum properties and gaining a better understanding of the quantum origin of their formation and their quantum noise limitations.
In this work, we experimentally study non-Hermitian lattice physics in photonic quadratures for the first time. Our quadrature lattice is formed by a tunable dissipative Kerr comb that couples different quadratures along the frequency dimension. We show theoretically and experimentally the relationship between dispersion symmetries and the squeezing properties of the multi-frequency lattice states of the below-threshold comb. Our work explores Kerr combs exhibiting multimode squeezed states with robust single peak frequency structure and double peak structure, thereby revealing fundamental properties related to the quantum nature of Kerr comb formation and quantum noise. We then observe such multimode states experimentally by measuring the anti-squeezing and photon statistics of the below-threshold Kerr comb. Our work demonstrates the relationship between non-Hermitian lattice physics and squeezed Kerr combs. To this end, we use the framework of non-Hermitian lattice physics and draw parallels with seminal classical non-Hermitian observations such as power oscillations caused by parity-time symmetry (PT symmetry) [15], and PT symmetric lattices [35, 36, 37]. This advancement will enable the generation of novel quantum optical resources on-chip, manipulation of quantum noise, and enhance the understanding of the quantum origin of Kerr comb formation.
@@ 2 Results
We consider a 2-FSR Kerr comb where every other cavity mode (every even-numbered cavity mode) is predominantly populated by coherent light, and light is pair-generated by the Kerr comb on all the other cavity modes (Fig. 1(a)). The loss in the odd cavity modes is greater than the parametric gain, such that these modes are populated by below-threshold light, i.e. a squeezed vacuum state. Each odd cavity mode is also directly coupled to neighboring odd modes (but not to the even modes) by Bragg scattering which originates predominantly from Kerr comb teeth separated in frequency by 2-FSR, creating a 1D lattice of frequency sites populated by vacuum states of light.
Figure 1: Squeezed quadrature lattices induced by Kerr combs. a. A 2-FSR Kerr comb (red) and a generated sub-threshold multimode state with squeezing frequency structure (green). b. General scheme of the system: A Kerr ring resonator is pumped by CW light, generating a 2-FSR bright comb (red circles). This comb generates a sub-threshold comb (green circles), which has the connectivity of a 1D lattice by pair generation and Bragg scattering. c. A unit cell of the 1D lattice in the bosonic creation and annihilation operator basis, longer range couplings decay with frequency distance.
Such a scenario can be evoked either by injecting a comb, as depicted in Fig. 1a into a Kerr resonator, or by pumping the resonator with continuous-wave (CW) light, causing the other comb teeth to emerge spontaneously via the symmetry-breaking phenomenon of Turing rolls [38] (Fig. 1(b)). The dynamics of the below-threshold light at the odd modes (centered around the pump at $\mu=0$) is given by the following Hamiltonian[7]

### v3

# Quadrature-Dependent Lattice Dynamics of Dissipative Microcombs

@@ Abstract
The study of coupled networks with parametric amplification of the vacuum fluctuations has garnered increasing interest due to its intricate physics and potential applications [1, 2, 3, 4, 5, 6]. In these systems, parametric interactions lead to beam-splitter coupling and two-mode squeezing, creating quadrature-dependent dynamics. These systems can be modeled as bosonic networks, arrays or lattices, exhibiting exotic effects such as unidirectional amplification and non-Hermitian chiral transport which influence multimode squeezing [7, 8, 9]. However, exploring and controlling these network dynamics experimentally in all-optical systems remains challenging. Recent advancements in integrated nonlinear micro-resonators, known as Kerr microcombs, enable the generation and control of broadband high-repetition pulses on microchips [10, 11, 12]. Kerr microcombs exhibit intriguing nonlinear dynamics, where coherent photons occupy discrete spectral lines, leading to multimode squeezed vacuum states [13, 14, 15]. In this work, we explore the lattice dynamics of vacuum fluctuations driven by dissipative Kerr microcombs. We design a photonic chip where a spontaneously emergent pair of pulses creates extended multimode states of parametrically amplified vacuum fluctuations. These states exhibit oscillatory dynamics, with implications on squeezing and secondary comb formation. By employing integrated micro-heaters, we tune the vacuum fluctuations to eliminate the oscillations, establishing a fundamental connection between non-Hermitian lattice symmetries and Kerr combs, paving the way for exotic quadrature-dependent optical networks with broad implications for quantum and classical photonic technologies.
@@ 1 Introduction
In recent years, there has been a growing interest in exploring the dynamics of multimode driven vacuum fluctuations [1, 9]. Traditionally, quantum vacuum fluctuations are parametrically amplified and subsequently undergo beam-splitting operations, leading to rich and potentially useful multimode squeezed states [16, 17, 18, 19, 20]. However, when parametric amplification and transport between different modes occur simultaneously, they induce dynamics of coupled networks or lattices, presenting a plethora of new effects from the perspective of squeezing and deterministic entanglement generation [3, 4, 1, 21, 22, 23]. These networks can be studied within the framework of the Bogoliubov modes and bosonic linear quadratic Hamiltonians. The eigenstates of the system, referred here as supermodes, are represented in the quadrature basis [24], which is also the natural basis for continuous-variable (CV) quantum states [25]. The quadratic bosonic Hamiltonians here exhibit non-trivial non-Hermiticity which is not from dissipation, as the underlying Hamiltonian is Hermitian. Nevertheless, the parametric processes give rise to non-Hermitian quadrature-dependent dynamics [2, 4, 26]. Such processes were recently explored in optomechanics within a synthetic frequency dimension [7, 8], exemplifying how symmetries and transport play a crucial role in the dynamics of the squeezed vacuum.
In the optical domain, multimode amplified quantum vacuum appears in both second-order and third-order nonlinear systems [27, 28, 29, 30]. In second-order nonlinear systems, photon transport between modes does not occur naturally and requires additional modulation to achieve lattice transport [6]. In contrast, third-order nonlinear processes present a different scenario. Remarkably, Kerr microcombs [10]—including dissipative Kerr solitons [11] and soliton crystals [12]—generated by third-order nonlinear resonators naturally produce both parametric amplification [31, 32] and photon transport between resonator modes through the Bragg scattering process [27, 29, 33, 34, 13]. This makes Kerr combs a unique system for studying multimode quantum phenomena [13, 35] and multimode squeezing [14, 15].
Consequently, Kerr combs can potentially induce different network geometries and lattice dynamics of multimode amplified vacuum fluctuations. This property of Kerr combs not only presents new methods for generating and controlling multimode quantum light but also enhances our understanding of classical nonlinear state evolution through the interplay between driven vacuum fluctuation dynamics and Kerr comb evolution in the resonator.
In this study, we investigate the lattice dynamics of optical vacuum fluctuations driven by dissipative Kerr micro-combs. We designed a microchip in which a pair of emerging pulses parametrically amplify and induce transport of the vacuum fluctuations, which subsequently exhibit lattice dynamics. We observe that the vacuum fluctuations generally oscillate, and that these oscillations are associated with a mismatch between the micro-comb repetition rate and the underlying dispersion. By utilizing integrated thermo-optic tunability, we are able to completely eliminate the oscillations and reach a stable, non-oscillatory (or stationary) regime of the amplified vacuum fluctuations. To explain this phenomenon, we draw parallels with classical parity-time (PT) symmetric lattices [36, 37, 38, 39]. Furthermore, we demonstrate how the oscillatory nature of the amplified vacuum fluctuations affects threshold processes, leading to the onset of (or lack of) RF beat notes and subsequent chaos in our micro-resonator. This advancement opens the door for modeling and utilizing bosonic network dynamics in an all-optical integrated platform with implication on multimode squeezing and entanglement between many frequency modes.
@@ 2 Results
To study the lattice dynamics of vacuum fluctuations driven by dissipative Kerr combs, we inject single-frequency (continuous wave, CW) light into a high-Q micro-resonator made of thin-film silicon carbide on insulator (more on the device in Methods 4.1) and spontaneously generate dissipative Kerr microcombs (Fig.1(a)). By red-tuning the input CW laser beam toward a resonance, a sequence of nonlinear transitions occurs, leading to the formation of different microcombs. Figure 1(b) shows the Lugiato–Lefever equations (LLE) simulation [40] of the power inside the resonator and corresponds to our device (see Methods 4.3 for more details on the simulation). The discontinuous steps in the transmission are nonlinear state transitions which correspond to different comb spectra (see Methods 4.2 for experimental comb spectra). For now, we focus on the first generated comb - the primary comb which appears after the first nonlinear threshold. Figure 1(c) illustrates that in our device this microcomb corresponds to the spontaneous formation of two pulses via the symmetry-breaking phenomenon of Turing rolls [41]. Accordingly, the spectrum of our two pulses forms a frequency comb with discrete frequency lines separated by twice the free spectral range (FSR), which is also the pulse repetition rate (blue lines in Fig.1(c)). In our system the FSR is 150 GHz.
To experimentally achieve the spontaneous formation of combs that include multiple frequency lines within a relatively narrow bandwidth (1 THz), we required strong anomalous dispersion. The dispersion of the material forming the waveguide is normal, and the sufficiently strong anomalous dispersion could not be achieved by geometric dispersion via tuning the waveguide cross section. Instead, we chose a multimode cross section and selected devices which feature an avoided mode crossing of two mode families - transverse electric (TE)/ transverse magnetic (TM). The mode crossing strongly detunes the resonance close to the pump frequency (Methods 4.1), which allows for an optical parametric oscillation (OPO) threshold between cavity modes that are separated in frequency by only 300 GHz, which is twice the FSR. This leads to the emergence of the 2-FSR primary Kerr comb, as illustrated in Fig. 1(c).
Figure 1: Scheme for exploring the lattice dynamics of vacuum fluctuations driven by dissipative Kerr microcombs. a. We inject a CW laser into a high-Q microring, which generate dissipative Kerr combs, and we measure the transmitted light. b. The nonlinearity and dispersion of our devices lead to several nonlinear transitions, as shown in the simulated comb power plot vs input laser frequency detuning, a solution of the Lugiato-Lefever equation for representative system parameters. Our main focus is the state (i) after threshold, in which two Turing roll pulses (forming a 2-FSR primary comb) spontaneously emerge, and the threshold after it leading to (ii) a secondary comb. Following thresholds lead to (iii) transition oscillations, and finally (iv) solitons. c. The spectrum of the two pulses shows that between each pair of neighboring comb lines (even modes), there is a resonance with amplified vacuum fluctuations (odd modes). The pulses induce transport (Bragg scattering), pair-generation (parametric amplification), and detuning (cross phase modulation) of the vacuum fluctuations. The lower panel shows how folding the frequency axis reveals lattice dynamics defined by a unit-cell of the vacuum fluctuations. The relatively narrow bandwidth of our pulses means that transport is governed by nearest-neighbor terms, which result in the lattice dynamics of the vacuum fluctuations. d. A simplified scheme of the lattice geometry of the lower panel of (c), each red circle represents a single unit-cell composed of two cavity modes
The choice of a 2-FSR primary Kerr comb with a relatively narrow bandwidth, confine the multimode dynamics and simplify the geometry of the driven vacuum fluctuations. The driven vacuum fluctuations that are the focus of our study are in the modes that are not populated with coherent light. To identify these modes we assign numbers $\mu$ to the resonances in our cavity. We define the pumped resonance as $\mu=0$ and thus the 2-FSR microcomb exists on every even-numbered cavity mode which are predominantly populated by coherent light. In contrast, light in the odd-numbered modes is dominated by vacuum fluctuations which are driven (pair generated) by the microcomb (indicated by red areas in Fig. 1(c)). This occurs as long as the cavity loss in the odd-numbered modes is greater than the parametric gain generated by the microcomb. The light remains in a driven vacuum state until the parametric gain increases above the cavity loss leading to a secondary threshold that generates more coherent comb lines. Next, we discuss the connectivity of the driven vacuum fluctuations and their modeling as a network of sites (cavity modes) and more specifically: a lattice. Figure 1(c,d) present the network model for our below threshold light in the odd-numbered modes, which has a lattice geometry. Transport, which is the result of Bragg scattering, connects primarily odd cavity modes to their neighboring odd cavity modes rather than to the even-numbered modes. This aspect makes our network into a one-dimensional lattice rather than a more general network of cavity modes populated by amplified vacuum states of light.
The lattice geometry is also related to the pair generation in our system. In Fig. 1(c) (lower panel), we plot the frequency axis in a “folded” form. In the folded form, modes $-\mu$ and $\mu$ are brought close together. The pair generation process is driven primarily by the pump mode $\mu=0$ together with the side-bands in cavity modes $\mu=\pm 2$ (see section 1 in the Supplementary Material). From energy conservation, pair generation occurs primarily between modes that are in proximity to one another within the folded geometry. Thus, if a photon is generated (or annihilated) in mode $\mu$ it’s pair will likely generate (or annihilate) at modes $-\mu,(-\mu+2)\text{ or }(-\mu-2)$. From similar considerations, Bragg scattering will be governed by transport of photons from mode $\mu$ to a close mode at $\mu=(\mu+2)\text{ or }(\mu-2)$. The simplified lattice model in the folded geometry is plotted in Fig.1(d). Taking into account all interactions, we can define a periodic unit cell consisting of four bosonic field operators: $a_{\mu}$, $a^{\dagger}_{\mu}$, $a_{-\mu}$, and $a^{\dagger}_{-\mu}$. The unit cell is the defining feature of the lattice model.
In what comes next, we solve the dynamics numerically, without relying on the simplified lattice model presented in Fig.1(c-d). The dynamics of the below-threshold light at the odd-numbered modes, centered around the pump at $\mu=0$, is governed by the following linearized Hamiltonian [13].